a) A monopoly market is characterized by the profit maximizer, price maker, high barriers to entry, single seller, and price discrimination. The information on the table below is a typical monopoly firm. Use it to answer the questions that follow


Output


Price


Total cost




$100


$200


1


100


290


2


100


350


3


100


390


4


100


470


5


100


560


6


100


700


7


100


900


i. What is the profit maximizing or loss minimizing quantity of output of this firm to produce? Why? 2marks


ii. What is the profit or loss at the profit maximizing or loss minimizing output level? 2marks


iii. Should the firm shut down or remain in operation? Why? 2marks


iv. Assuming this is a representative firm in the industry or market, given the profits or losses incurred, what can be expected to happen over time in the market? 2marks


1
Expert's answer
2021-05-30T15:31:22-0400
  1. Marginal revenue (MR) = the total revenue generated by increasing the output by an additional one unit.

Total Revenue (TR) = Price"\times" quantity

the marginal revenue and total revenue are found as shown in the following table:




For the firm to maximize profits, marginal revenue must be greater than marginal cost. If the monopoly firm produces a lower quantity, the MR is greater than MC at those levels of output and in this case the firm can only make higher profits by expanding output.


The profit maximizing quantity of output of this firm is at output =6 where the marginal cost rises above the marginal revenue.


(2) The profit at the profit maximizing output level =MC - MR

= 140 - 100

=40

(3 )The firm should remain in operation in order to increase the level of output so as to achieve the goal of maximizing profits.


(4) According to the profits that accrue as from output level =6 onwards, it is expected that over time, the monopoly is likely to capture a greater market and make more profits on the continual increase in level of output.